You have a sample of gas in a flask with a volume of 250 mL. At 25.5 °C, the pressure

of the gas is 360 mm Hg. If you decrease the temperature to −5.0 °C, what is the gas

pressure at the lower temperature?


1
Expert's answer
2021-10-01T13:11:01-0400

"\frac{P_1}{T_1}=\frac{P_2}{T_2}"


"\frac{360}{(273+25.5)}=\frac{P_2}{(273-5)}"


"P_2=\frac{360×268}{298.5}=323.22mmHg"